OpenStudy (anonymous):

Find the derivative of f(x) = 6x + 2 at x = 1.

OpenStudy (anonymous):

derivative of 6x + 2 is 6

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

the derivative of a.x^n is a.n.x^(n - 1) and the derivative of a constant is 0 so, the derivative of 6x + 2 = 6.1.x^(1-1) + 0 = 6.1.x^0 = 6.1 = 6